Check my work QS 1-8 Applying the accounting equation LO A1 1. Use the accounting equation to compute the missing financial statement amounts. 41,000 26,000| +|S 72,000 77,000 =IS $ 89,00022,000 2. Use the expanded accounting equation to compute the missing financial statement amounts AssetsLiabilities + $ 44,00017,00022,000-S S 83,800 S 33,000 46,000 ny -Dividends+RevenuesExpenses 9,000 26,000 19.000 < Prev 20, 12 Next > sling
Answer #1

Accounting Equation

Assets = liabilities + Equity

Company 1 ) Liabilites = Assets - Equity ( 77000-41000) = 36000

Company 2) Assets = labilites+ Equity ( 26000+72000) = 98,000

Company 3) Equty = assets - Laibilities ( 89000-22000) = 67000

Accounting Equation Can be Expanded

Assets = Liabilites +common Stock- Dividend +Revenue -Expense

Companv |Assets Liabilites Equit 1 77000 2 98000 3 89000 36000+ 26000+ 22000+ 41000 72000 67000 Liabilites+Common Stock 22000 46000 Dividend + Revenues Expenses 9000 19000 Companv |Assets 1 44000 2 83800 17000+ 33000+ +14000 26000 2200+